find volume of solid generated by revolving the region bounded by y = √(x) and the lines y = 2 and x = 0 about the x-axis
1
Expert's answer
2020-12-14T15:22:17-0500

Let us find volume of solid generated by revolving the region bounded by "y = \sqrt x" and the lines "y = 2" and "x = 0" about the "x" -axis.


Let us sketch the graph of the region bounded by "y = \sqrt x" and the lines "y = 2" and "x = 0":




"V=\pi\int\limits_0^4(2^2-(\sqrt x)^2)dx=\pi\int\limits_0^4(4- x)dx=\pi(4x-\frac{x^2}{2})|_0^4=\pi(16-\frac{16}{2})=8\pi"